Operation Summary for September 8, 2010 - September 10, 2010 |
Wednesday Sept. 8 The day shift began with Operations monitoring the Tevatron (TeV) store 8064 (91.12E30) and the Recycler stash (317.4E10), with the Antiproton Source (Pbar) stacking (22.5E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments off in preparation for accesses, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations terminated store 8064 at 10:56 AM, put the TeV into HEP shot setup, and established store 8065 at 1:38 PM with an average initial luminosity (AIL) of 128.35E30. Operations resumed sending beam to NuMI for target scans at 3:15 PM. During the evening shift, Operations terminated store 8065 at 4:34 PM. Operations began sending beam to MiniBooNE at 5:20 PM. Operators prepared CDF for an access at 6:08 PM that lasted until 10:54 PM. Operations put the TeV through a dry squeeze. Thursday Sept. 9 The midnight shift began with operators conducting a TeV orbit check, with Operations monitoring the stash (327E10), with Pbar stacking (44.8E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations put the TeV into shot setup, but this was cut short due to TCLK errors on a couple of TeV corrector magnet power supply ramp cards. Operators reset the cards, conducted another orbit check, and then put the TeV into shot setup at 3:11 AM. Operations established store 8072 at 4:11 AM with an AIL of 262.3E30. During the day shift, Operations terminated store 8072 at 2:30 PM, put the TeV into shot setup, and established store 8073 at 3:50 PM with an AIL of 250.4E30. During the evening shift, Operations reported that BPM issues kept the NuMI beam off until 5:26 PM. NuMI resumed taking beam until a problem with the ‘Baratron Helium Pressure’ sensor at 9:43 PM required an access into the NuMI target hall. Later, NuMI personnel determined that there was a helium leak in the target and MI personnel determined that a MI kicker (KPS1C) thyratron tube was failing. Friday Sept. 10 The midnight shift began with Operations monitoring store 8073 (91.04E30) and the stash (183.2E10), with Pbar stacking (26.9E10), with NuMI personnel investigating their helium leak, and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8073 at 3:05 AM. Operators prepared D0 for an access. After the access ended, Operations conducted a TeV checkout and then put the TeV into shot setup. Operations began sending NuMI a low intensity beam for target scans at 3:20 AM. MI resumed slip stacking and sending normal beam to NuMI at 5:23 AM. Operations established store 8077 at 7:24 AM with an AIL of 248.1E30. Operations halted Booster beam at 7:31 AM for an access later in the morning. The Plans for Friday and the Weekend The plans for today are to work on the NuMI helium leak, the MI kicker Thyratron, a Booster RF station (BRF5) repair, and a Linac RF station (LRF1) power supply repair. The plans for the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 09/10/2010 email Fermilab |
| Security, Privacy, Legal | |